# Humanoid World Models [Uncaptioned image] : Open World Foundation Models for Humanoid Robotics

### v1

@@ Abstract
Humanoid robots have the potential to perform complex tasks in human-centered environments but require robust predictive models to reason about the outcomes of their actions. We introduce Humanoid World Models (HWM), a family of lightweight, open-source video-based models that forecast future egocentric observations conditioned on actions. We train two types of generative models—Masked Transformers and Flow-Matching—on 100 hours of humanoid demonstrations. Additionally, we explore architectural variants with different attention mechanisms and parameter-sharing strategies. Our parameter-sharing techniques reduce model size by 33–53% with minimal impact on performance or visual fidelity. HWM is designed to be trained and deployed in practical academic and small-lab settings, such as 1–2 GPUs.
@@ 1 Introduction
Autonomous humanoid robots have the potential to transform both industry and daily life by automating tasks that are dull, dangerous, or physically demanding (Goswami & Vadakkepat, 2019). Their human-like morphology allows them to operate in spaces built for people, interact naturally with humans, and easily learn from teleoperated demonstrations. Realizing this promise, however, requires more than hardware: these agents must be equipped with sophisticated models of reasoning to navigate the complexity and unpredictability of real-world environments.
@@
While large multimodal models (Bommasani et al., 2022; Zhang et al., 2024; Kim et al., 2024; Black et al., 2024) show some promise in generalization and reasoning, they often lack the accuracy, reliability, and robustness needed for embodied AI in open-world settings (Tong et al., 2024; Pfeifer & Iida, 2004). As a result, they struggle to meet the demands of embodied agents that must act safely and effectively in dynamic, unstructured environments (Pfeifer & Iida, 2004; Li et al., 2024; Duan et al., 2024).
Figure 1: Overview of Humanoid World Model.
@@
One avenue for improving humanoid intelligence and control is through World Models—predictive models trained to forecast future outcomes based on past observations and actions (Ha & Schmidhuber, 2018). In our setting, these function as action-conditioned video generators: they predict future visual states as sequences of frames, enabling agents to simulate the consequences of candidate actions (Du et al., 2023). By simulating the outcomes of actions before trying them in the real world, world models support long-horizon planning, counterfactual reasoning, and data-efficient policy learning through synthetic rollouts (Yang et al., 2024, 2023).
@@
Despite recent progress in video generation, most models are built for entertainment applications (Liu et al., 2024; Polyak et al., 2024), emphasizing visual appeal over physical, ego-centric plausibility. Many world models remain closed-source, require large-scale compute for training or inference, or are not designed for humanoid robots (Yang et al., 2023; NVIDIA et al., 2025; Zhu et al., 2024). As a result, there is a clear gap: few open-source models are both physically grounded for humanoid robots and lightweight enough to train or run inference on modest academic hardware (e.g., 2–3 GPUs). To address this, we ask: Can we build a physically plausible, humanoid-specific world model that runs on just two GPUs?
@@
We introduce Humanoid World Models (HWM), an open-source lightweight world model for humanoid robotics on 100 hours of humanoid video demonstrations. We investigate two distinct video generation paradigms: Masked Transformers and Flow-Matching models. Drawing from recent advances in image and video generation, we explore four architectural variations within each framework. These variations are defined along two axes: (1) joint vs. cross-attention mechanisms, and (2) shared vs. separate parameters across token streams. This design space is motivated by successful architecture strategies from the much more explored text-to-image models.
@@
Our experiments show that, given our dataset and compute constraints, Masked Transformers consistently outperformed Flow-Matching models—even when the latter used more parameters and were trained for longer. Across both families of models, we observed that the different architectural variants performed comparably in most scenarios. However, key trends emerged: within the Masked Transformer framework, the joint attention variant achieved the best overall performance. In contrast, for Flow-Matching models, split attention proved most effective.
@@ 1.2 World Models
A core challenge in robotics is enabling agents to perceive their environment, reason over possible actions, and plan goal-directed behavior—especially in novel or unstructured settings. Foundation models (Bommasani et al., 2022) attempt to generalize across tasks through large-scale, multimodal training. Recent approaches include Vision-Language-Action (VLA) models that predict actions from video and language inputs (Black et al., 2024; Kim et al., 2024), and large language or vision-language models used as high-level planners (Liang et al., 2023; Li et al., 2024; Wang et al., 2024). However, these methods often struggle with spatial reasoning (Tong et al., 2024), continuous sensorimotor processing, and require complex prompting strategies (Li et al., 2024).
@@ 1.3 Video Generation Models
Large-scale text-to-video diffusion models such as Sora (Liu et al., 2024), CogVideoX (Yang et al., 2025), and MovieGen (Polyak et al., 2024) achieve impressive visual quality but are designed for entertainment and lack support for conditioning on past video—a key requirement for physically grounded, ego-centric prediction.
@@ 2.1 Formulation
Following prior works (Rombach et al., 2022; Chang et al., 2022a), we train our generative models in a VAE’s compressed latent space: $v_{p}$ and $v_{f}$ are encoded into latent representations $L_{p}$ and $L_{f}$. Masked-HWM uses VQ-VAE to quantize latents into tokens from a finite vocabulary of size $s$, while Flow-HWM uses a continuous VAE.
@@ 2.2 Masked Video Modelling
We train the Masked-HWM variant using the Masked Video Modelling (MVM) paradigm (Chang et al., 2022a; Yu et al., 2023). After passing passing $v_{p},v_{f}$ through a VQ-VAE to yield $\mathbf{L_{p}},\mathbf{L_{f}}$, we concatenate the past and future latent tokens $\mathbf{L}=[\mathbf{L_{p}};\mathbf{L_{f}}]$ along the temporal dimension.
@@ 2.4 Transformer Block Design
The Base Block design for the Masked-HWM is illustrated in Figure 2. Following prior work in non-autoregressive video generation methods (Bruce et al., 2024; Xiang et al., 2024), we adopt a factorized or separate spatial and temporal attention layers. Compared to full spatiotemporal attention, factorized attention reduces the computational cost and scales more efficiently with video length. During temporal attention, all tokens from various streams $[a_{p},a_{f},\mathbf{X}]$ jointly attend to one another along the temporal dimension. Spatial attention is applied seperately only to the video tokens. Rotary Position Embeddings (RoPE)(Su et al., 2023) are used during spatial and temporal attention.
@@ 3.2.2 Quantitative Results
Overall, none of the Flow-HWM variants outperformed the Masked-HWM models in either visual quality or sampling speed, suggesting that masked video modeling is a more effective generative paradigm for our dataset and compute constraints. The results consistently show that parameter sharing—particularly in the Full Sharing variant—provides substantial efficiency gains with minimal loss in visual fidelity, very benefitial in resource-constrained settings.

### v2

@@ Abstract
Humanoid robots, with their human-like form, are uniquely suited for interacting in environments built for people. However, enabling humanoids to reason, plan, and act in complex open-world settings remains a challenge. World models, models that predict the future outcome of a given action, can support these capabilities by serving as a dynamics model in long-horizon planning and generating synthetic data for policy learning. We introduce Humanoid World Models (HWM), a family of lightweight, open-source models that forecast future egocentric video conditioned on humanoid control tokens. We train two types of generative models, Masked Transformers and Flow-Matching, on 100 hours of humanoid demonstrations. Additionally, we explore architectural variants with different attention mechanisms and parameter-sharing strategies. Our parameter-sharing techniques reduce model size by 33–53% with minimal impact on performance or visual fidelity. HWMs are designed to be trained and deployed in practical academic and small-lab settings, such as 1–2 GPUs.
@@ 1 Introduction
Autonomous humanoid robots have the potential to transform both industry and daily life by automating tasks that are dull, dangerous, or physically demanding (Goswami & Vadakkepat, 2019). Their human-like morphology allows them to operate in spaces built for people, interact naturally with humans, and easily learn from teleoperated demonstrations. However, in order to navigate the complexity and unpredictability of real-world environments these agents require sophisticated reasoning capabilities.
@@
While large multimodal models (Zhang et al., 2024; Black et al., 2024) show some promise in reasoning tasks, they often lack the accuracy, reliability, and robustness needed for embodied AI in open-world settings (Tong et al., 2024; Pfeifer & Iida, 2004). As a result, they struggle to meet the demands of embodied agents that must act safely and effectively in dynamic, unstructured environments (Pfeifer & Iida, 2004; Li et al., 2024; Duan et al., 2024).
Figure 1: Overview of Humanoid World Models (HWM). Given past video observations and humanoid control tokens (joint angles, velocities, etc.), it predicts future video observations.
@@
One avenue for improving humanoid intelligence and control is through World Models—predictive models trained to forecast future outcomes based on past observations and actions (Ha & Schmidhuber, 2018). In our setting, these function as action-conditioned video generators: they predict future visual states as sequences of frames, enabling agents to simulate the consequences of candidate actions (Du et al., 2023). By simulating the outcomes of actions without trying them in the real world, world models can be used for long-horizon planning by using them as a dynamics model (Yang et al., 2024) and enable more data-efficient policy learning through synthetic rollouts (Yang et al., 2023).
@@
Despite recent progress in video generation, most models are built for entertainment applications (Liu et al., 2024; Polyak et al., 2024), emphasizing visual appeal over physical, ego-centric plausibility. Many world models remain closed-source, require large-scale compute for training or inference, or are not designed for humanoid robots (Yang et al., 2023; NVIDIA et al., 2025; Zhu et al., 2024). As a result, there is a clear gap: few open-source models are both physically grounded for humanoid robots and lightweight enough to train and deploy on modest academic hardware (e.g., 2–3 GPUs). To address this, we ask: Can we build a physically plausible, humanoid-specific world model that can be trained and deployed on as little as two GPUs?
@@
We introduce Humanoid World Models (HWM), a set of open-source, lightweight world models for humanoid robotics trained on 100 hours of humanoid video demonstrations. We investigate two distinct video generation paradigms: Masked Transformers and Flow-Matching models. Drawing from recent advances in image and video generation, we explore four architectural variations within each framework. These variations are defined along two axes: (1) joint vs. cross-attention mechanisms, and (2) shared vs. separate parameters across token streams. This design space is motivated by successful architecture strategies from the more well-studied text-to-image model literature.
@@
Our experiments show that, given our dataset and compute constraints, Masked Transformers consistently outperformed Flow-Matching models even when the latter used more parameters and were trained for longer. Across both families of models, we observed that the different architectural variants performed comparably in most scenarios. However, key trends emerged: within the Masked Transformer framework, the joint attention variant achieved the best overall performance. In contrast, for Flow-Matching models, split attention proved most effective.
@@ 1.2 World Models
A core challenge in robotics is enabling agents to perceive their environment, reason over possible actions, and plan goal-directed behavior, especially in novel or unstructured settings. Foundation models (Bommasani et al., 2022) attempt to generalize across tasks through large-scale, multimodal training. Recent approaches include Vision-Language-Action (VLA) models that predict actions from video and language inputs (Black et al., 2024; Kim et al., 2024), and large language or vision-language models used as high-level planners (Liang et al., 2023; Li et al., 2024; Wang et al., 2024). However, these methods often struggle with spatial reasoning (Tong et al., 2024), continuous sensorimotor processing, and require complex prompting strategies (Li et al., 2024).
@@ 1.3 Video Generation Models
Large-scale text-to-video diffusion models such as Sora (Liu et al., 2024), CogVideoX (Yang et al., 2025), and MovieGen (Polyak et al., 2024) achieve impressive visual quality but are designed for entertainment and lack support for conditioning on past video, a key requirement for physically grounded, ego-centric prediction.
@@ 2.1 Formulation
Following prior works (Rombach et al., 2022; Chang et al., 2022a), we train our generative models in a VAE’s compressed latent space: $v_{p}$ and $v_{f}$ are encoded into latent representations $L_{p}$ and $L_{f}$. Masked-HWM uses a VQ-VAE to quantize latents into tokens from a finite vocabulary of size $s$, while Flow-HWM uses a continuous VAE.
@@ 2.2 Masked Video Modelling
We train the Masked-HWM variant using the Masked Video Modelling (MVM) paradigm, inspired by MaskGIT and MAGVIT (Chang et al., 2022a; Yu et al., 2023). After passing passing $v_{p},v_{f}$ through a VQ-VAE to yield $\mathbf{L_{p}},\mathbf{L_{f}}$, we concatenate the past and future latent tokens $\mathbf{L}=[\mathbf{L_{p}};\mathbf{L_{f}}]$ along the temporal dimension.
@@ 2.4 Transformer Block Design
The Base Block design for the Masked-HWM is illustrated in Figure 2. Following prior work in non-autoregressive video generation methods (Bruce et al., 2024; Xiang et al., 2024), we adopt a factorized or separate spatial and temporal attention layers. Compared to full spatiotemporal attention, factorized attention reduces the computational cost and scales more efficiently with video length. During temporal attention, all tokens from various streams $[a_{p},a_{f},\mathbf{L}]$ jointly attend to one another along the temporal dimension. Spatial attention is applied seperately only to the video tokens. Rotary Position Embeddings (RoPE)(Su et al., 2023) are used during spatial and temporal attention.
@@ 3.2.2 Quantitative Results
Overall, none of the Flow-HWM variants outperformed the Masked-HWM models in either visual quality or sampling speed, suggesting that masked video modeling is a more effective generative paradigm for our dataset and compute constraints. The results consistently show that parameter sharing, particularly in the Full Sharing variant, provides substantial efficiency gains with minimal loss in visual fidelity, very benefitial in resource-constrained settings.